# A Retrieval Augmented Spatio-Temporal Framework for Traffic Prediction

### v1

@@ Abstract
Traffic prediction is a cornerstone of modern intelligent transportation systems and a critical task in spatio-temporal forecasting. Although advanced Spatio-temporal Graph Neural Networks (STGNNs) and pre-trained models have achieved significant progress in traffic prediction, two key challenges remain: (i) limited contextual capacity when modeling complex spatio-temporal dependencies, and (ii) low predictability at fine-grained spatio-temporal points due to heterogeneous patterns.
Inspired by Retrieval-Augmented Generation (RAG), we propose RAST, a universal framework that integrates retrieval-augmented mechanisms with spatio-temporal modeling to address these challenges. Our framework consists of three key designs: 1) Decoupled Encoder and Query Generator to capture decoupled spatial and temporal features and construct a fusion query via residual fusion; 2) Spatio-temporal Retrieval Store and Retrievers to maintain and retrieve vectorized fine-grained patterns; and 3) Universal Backbone Predictor that flexibly accommodates pre-trained STGNNs or simple MLP predictors. Extensive experiments on six real-world traffic networks, including large-scale datasets, demonstrate that RAST achieves superior performance while maintaining computational efficiency.
I Introduction
Traffic prediction stands as a cornerstone of modern Intelligent Transportation Systems (ITS), enabling critical applications including traffic management, route optimization, and congestion mitigation [1, 2]. The accurate forecasting of traffic conditions directly impacts urban mobility, economic efficiency, and environmental sustainability across metropolitan areas worldwide. Spatio-temporal Forecasting (STF) provides the methodological foundation for addressing these traffic prediction challenges, as traffic data inherently exhibits complex interdependencies across both spatial and temporal dimensions [3, 4].
@@ I Introduction
The evolution of STF methodologies has progressed from traditional statistical approaches [5, 6, 7] to sophisticated deep learning architectures [8, 9, 10] as application scenarios have become increasingly complex. This progression culminated in the development of Spatio-temporal Graph Neural Networks (STGNNs) [11, 12, 13], which have achieved remarkable success in modeling complex spatial-temporal dependencies by representing traffic networks as graphs and leveraging graph convolution operations [14, 15]. More recently, the emergence of Large Models (LMs) and pre-trained models from Natural Language Processing (NLP) [16, 17] and Computer Vision (CV) [18, 19] has opened new opportunities for enhancing spatio-temporal forecasting capabilities [20, 21, 22].
Despite these advances, current STF approaches face two critical limitations: (i) Limited Contextual Capacity vs. scale of ST data: Contemporary pre-trained STGNNs suffer from constrained contextual embedding capacity when handling complex spatio-temporal dependencies in large-scale traffic networks [23, 24, 25]. Drawing inspiration from retrieval-augmented generation (RAG) that has shown promise in addressing context limitations in Large Language Models (LLMs) [26], we investigate whether retrieval-augmented mechanisms can compensate for the limited spatio-temporal learning capacity, as illustrated in Figure 1; (ii) Complex Architecture vs. Low Predictability: Due to the inherent heterogeneity in spatio-temporal data [4], existing STF approaches lack efficient mechanisms for fine-grained pattern adjustment within limited embedding lengths [4, 23, 24]. Current performance improvements of the STGNNs rely heavily on complex model architectures [27, 28] to capture overall trends, yet low-predictability points in both temporal and spatial dimensions remain difficult to capture. Instead of further increasing model complexity by adding more weighted parameters, we propose to capture complex spatio-temporal dependencies through explicit memory storage and retrieval mechanisms, as demonstrated in Figure 2.
@@
To bridge the gap in applying RAG to STF and further address the aforementioned challenges, we propose RAST (Retrieval-Augmented Spatio-Temporal forecasting), a universal framework that integrates retrieval-augmented mechanisms with spatio-temporal modeling for traffic prediction and pre-trained model enhancements. Specifically, our approach maintains a vector-based dual-dimension spatio-temporal retrieval store. During training, our method decouples input data into spatial and temporal encodings used to update the store while constructing context-aware queries through residual fusion. ST-Retrievers then search and project dual-dimension retrieval embeddings under queries. Finally, we fuse decoupled retrieval embeddings with current queries through a cross-attention module and obtain predictions via the universal backbone predictor. Our contributions are summarized as follows:
@@
  • Comprehensive Empirical Validation: Extensive experiments on six real-world datasets demonstrate that our proposed method effectively captures complex spatio-temporal patterns while maintaining high computational efficiency, achieving up to 24.75% improvement in average MAE compared to the RPMixer on the SD dataset.
@@ II Related Work
is a fundamental task in numerous application domains such as traffic management, urban planning, and environmental monitoring [29, 4, 30, 31, 32]. Early deep learning methods combined CNNs and RNNs [8, 10] but struggled with non-Euclidean traffic networks. The emergence of Spatio-temporal Graph Neural Networks (STGNNs) addressed this limitation by integrating Graph Neural Networks (GNNs) [9] with temporal models [14, 33, 12, 13, 34, 27, 35, 11, 15, 36]. Despite these advancements, the performance improvements of STGNNs have begun to plateau due to limited contextual capacity and reliance on increasingly complex architectures [37, 23]. This stagnation has prompted research toward integrating pre-trained models and Large Language Models (LLMs) [20, 38, 21, 39, 40] to enhance predictive capabilities. However, these methods still struggle to capture spatio-temporal heterogeneity in large-scale scenarios, while universal solutions for contextual capacity limitations in spatio-temporal modeling remain largely unexplored.
@@
has emerged as a transformative paradigm for enhancing large language model performance, particularly in knowledge-intensive tasks [41, 26, 42]. Traditional LLMs struggle with tasks requiring vast amounts of factual knowledge or domain-specific expertise due to finite parametric memory and limited context [43, 44]. RAG addresses this limitation by integrating external retrieval mechanisms that enable models to dynamically access relevant information from large knowledge bases during inference [45, 46, 47], improving open-domain question answering, fact-checking, and few-shot learning tasks [48, 49]. While the retrieval-augmented mechanism has been extensively studied in natural language processing, its application to vast spatio-scenarios presents significant opportunities, particularly given the similar challenges of constrained model capacity.
III Methodology
In this section, we present RAST, a Retrieval Augmented Spatio-Temporal forecasting framework as illustrated in Figure 3.
Our approach consists of three core components working synergistically: 1) Decoupled Encoder Layers that convert raw inputs into vectorized embedding representations, 2) Query Generator that constructs fusion queries via residual fusion, 3) Retrieval Store that maintains and fine-grained historical patterns 4) ST-Retriever that retrieves embeddings based on queries and further fuses them with cross-attention blocks, and 5) Backbone Predictor that accommodates diverse pre-trained STGNNs or simple predictors (e.g. MLP) for enhanced forecasting performance.
III-A Problem Formulation
@@
III-B Data Encoding and Query Construction
@@
Following recent advances in spatio-temporal modeling [50], we employ decoupled encoder layers to separately process temporal and spatial information. This encoding module initially captures basic characteristics (e.g., cyclicity for temporal, regionality for spatial) formulated as follows:
@@
$\mathbf{E}^{(l+1)}$ $=\text{LayerNorm}(\mathbf{E}_{f}^{(l)}+\text{FFN}(\mathbf{E}_{f}^{(l)})),$ (5)
@@
III-C Spatio-temporal Retrieval Store
@@
$\mathcal{M}_{sp}=\{\mathbf{v}^{(i)}_{sp},\mathbf{m}_{sp}^{(i)}\}_{i=1}^{|\mathcal{M}_{sp}|},\mathcal{M}_{tp}^{(j)}=\{\mathbf{v}_{tp}^{(j)},\mathbf{m}_{tp}^{(j)}\}_{j=1}^{|\mathcal{M}_{tp}|}$ (7)
@@
To enable fast similarity search, our retrieval store utilizes the Facebook AI Similarity Search (FAISS) library [51, 52] for efficient similarity-based indexing. Given history queries $\mathcal{Q}_{st}$ and current state $e$, we maintain and compute indices $\mathcal{I}$ for temporal and spatial embeddings (more details in Appendix D):
@@
III-C1 Information-Theoretic ST-Retriever
@@
$\mathcal{E}_{s}$ $=\text{Retriever}(\mathbf{E}_{sp},\mathcal{I}_{s},k)=\{(\mathbf{v}^{(i)}_{s},\omega_{s}^{(i)}\}_{i=1}^{k}$ (12)
@@
III-C2 Momentum-Based Memory Management
@@ III-D Prediction and Optimization
RAST (Ours) 15.36 25.81 16.47 18.39 29.93 12.43 19.52 32.73 8.23 14.20 23.49 9.29
TABLE I: Performance comparison assessed by averaging over all 12 prediction steps with baseline models on the PEMS03, 04, 07, 08 datasets. Bold: best; Underline: second best.
IV Experiments
@@ IV-A Experimental Settings
We conducted comprehensive experiments across six diverse traffic networks: 1) PEMS03, PEMS04, PEMS07, PEMS08 [53] datasets, and 2) the large-scale dataset San Diego (SD), Greater Bay Area (GBA) as introduced in LargeST [25]. Detailed statistics of these datasets are shown in Table II. We compare our RAST with 21 classic or advanced baselines, which are categorized into 3 groups: (i) Non-spatial methods: ARIMA [5]VAR [6], SVR [54], LSTM [10], TCN [55], Transformer [56]; (ii) GNN-based Spatial-temporal Models DCRNN [14], STGCN [12], ASTGCN [34], GWNet [15], LSGCN [57], STSGCN [53], STFGNN [58], STGODE [59], DSTAGNN [28], AGCRN [13], D2STGNN [27] and (iii) Other Enhanced Approaches: Z-GCNETs [60], TAMP [61], STKD [62], RPMixer [63]. More details of the baselines are given in Appendix A.
@@
Performance is evaluated using standard metrics, including MAE, RMSE, and MAPE. We use 12 historical time steps to forecast the next 12 steps and calculate the average across horizons 3, 6, and 12. For the fairness of the experiment, we only utilized the simple MLP trained from scratch as the backbone predictor in the following experiments. All models are implemented in PyTorch and trained on Nvidia RTX A6000 GPUs with consistent hyperparameter tuning protocols to ensure fair comparison. The models are trained using the Adam optimizer with a learning rate of 0.002, a batch size of 32, and a maximum of 300 epochs, applying an early stopping strategy. More detailed experimental settings are provided in Appendix B.
Datasets #Points #Samples #TimeSlices Timespan
PEMS03 358 9.38M 26,208 09/01/2018-11/30/2018
PEMS04 307 5.22M 16,992 01/01/2018-02/28/2018
PEMS07 883 24.92M 28,224 05/01/2017-08/31/2017
PEMS08 170 3.04M 17,856 07/01/2016-08/31/2016
SD 716 25M 35040 01/01/2019-12/31/2019
GBA 2352 82M 35040 01/01/2019-12/31/2019
TABLE II: The Statistics Details of the Dataset.
Datase Methods Horizon 3 Horizon 6 Horizon 12 Average
@@
RAST (Ours) 16.23 26.75 10.23 19.12 32.05 12.45 23.20 40.72 16.06 19.00 32.64 12.53
GBA
LSTM 20.38 33.34 15.47 27.56 43.57 23.52 39.03 60.59 37.48 27.96 44.21 24.48
@@
RAST (Ours) 18.57 30.12 14.60 22.12 35.54 18.06 26.31 43.14 21.75 21.75 35.82 17.90
TABLE III: Large-scale traffic forecasting performance comparison of our RAST and baselines on the SD and GBA datasets. Our RAST achieves the best performance across all prediction horizons and metrics. Bold: best; Underline: second best.
IV-B Performance Evaluation (RQ1)
Table I presents a comprehensive comparison of various approaches for traffic forecasting tasks across PEMS datasets. While the results on PEMS03 show room for improvement, RAST consistently outperforms state-of-the-art baseline methods on the remaining datasets. On the PEMS07 dataset, RAST achieves a MAE of 19.52, surpassing the second-best method DSTAGNN by 8.87%. On the PEMS08 dataset, RAST reduces RMSE by 1.58 compared to competitive STKD, while on the PEMS04, it outperforms all baselines with an MAE of 18.39 and RMSE of 29.93.
Table III extends our evaluation to larger datasets, while the performance on the horizon 3, horizon 6, horizon 12, and the average of the whole 12 horizons are reported. Our method maintains its advantage on larger traffic networks, achieving the best results at each horizon. On the SD dataset, RAST achieves an average MAE of 19.00, RMSE of 32.64, and MAPE of 12.53%, representing improvements compared to the second-best method STGODE. For the long-term prediction (Horizon 12), our method achieves the best MAE of 23.20 and RMSE of 40.72, highlighting its robustness in modeling complex long-range dependencies. On even larger traffic datasets GBA, RAST still surpasses baselines across all metrics and horizons, outperforming strong baselines like DSTAGNN and RPMixer. These results demonstrate the scalability of RAST to model complex spatio-temporal dependencies in large-scale traffic networks.

### v2

# RAST: A Retrieval Augmented Spatio-Temporal Framework for Traffic Prediction

@@ Abstract
Traffic prediction is a cornerstone of modern intelligent transportation systems and a critical task in spatio-temporal forecasting. Although advanced Spatio-temporal Graph Neural Networks (STGNNs) and pre-trained models have achieved significant progress in traffic prediction, two key challenges remain: (i) limited contextual capacity when modeling complex spatio-temporal dependencies, and (ii) low predictability at fine-grained spatio-temporal points due to heterogeneous patterns. Inspired by Retrieval-Augmented Generation (RAG), we propose RAST, a universal framework that integrates retrieval-augmented mechanisms with spatio-temporal modeling to address these challenges. Our framework consists of three key designs: 1) Decoupled Encoder and Query Generator to capture decoupled spatial and temporal features and construct a fusion query via residual fusion; 2) Spatio-temporal Retrieval Store and Retrievers to maintain and retrieve vectorized fine-grained patterns; and 3) Universal Backbone Predictor that flexibly accommodates pre-trained STGNNs or simple MLP predictors. Extensive experiments on six real-world traffic networks, including large-scale datasets, demonstrate that RAST achieves superior performance while maintaining computational efficiency.
Codehttps://github.com/RWLinno/RAST
1 Introduction
Traffic prediction stands as a cornerstone of modern Intelligent Transportation Systems (ITS), enabling critical applications including traffic management, route optimization, and congestion mitigation (chavhan2020prediction; zheng2014urban). The accurate forecasting of traffic conditions directly impacts urban mobility, economic efficiency, and environmental sustainability across metropolitan areas worldwide. Spatio-temporal Forecasting (STF) provides the methodological foundation for addressing these traffic prediction challenges, as traffic data inherently exhibits complex interdependencies across both spatial and temporal dimensions (wang2020deep; jin2023spatio).
@@ 1 Introduction
The evolution of STF methodologies has progressed from traditional statistical approaches (box2015time; lutkepohl2005new; chandra2009predictions) to sophisticated deep learning architectures (shi2015convolutional; kipf2016semi; hochreiter1997long) as application scenarios have become increasingly complex. This progression culminated in the development of Spatio-temporal Graph Neural Networks (STGNNs) (wu2020connecting; yu2017spatio; bai2020adaptive), which have achieved remarkable success in modeling complex spatial-temporal dependencies by representing traffic networks as graphs and leveraging graph convolution operations (li2017diffusion; wu2019graph). More recently, the emergence of Large Models (LMs) and pre-trained models from Natural Language Processing (NLP) (devlin2018bert; jin2021trafficbert) and Computer Vision (CV) (dosovitskiy2020image; he2022masked) has opened new opportunities for enhancing spatio-temporal forecasting capabilities (zhou2024one; jin2024position; yan2024urbanclip).
Despite these advances, current STF approaches face two critical limitations: (i) Limited Contextual Capacity vs. scale of ST data: Contemporary pre-trained STGNNs suffer from constrained contextual embedding capacity when handling complex spatio-temporal dependencies in large-scale traffic networks (jin2024survey; jiang2023deep; liu2023largest). Drawing inspiration from retrieval-augmented generation (RAG) that has shown promise in addressing context limitations in Large Language Models (LLMs) (lewis2020retrieval), we investigate whether retrieval-augmented mechanisms can compensate for the limited spatio-temporal learning capacity, as illustrated in Figure 1; (ii) Complex Architecture vs. Low Predictability: Due to the inherent heterogeneity in spatio-temporal data (jin2023spatio), existing STF approaches lack efficient mechanisms for fine-grained pattern adjustment within limited embedding lengths (jin2023spatio; jin2024survey; jiang2023deep). Current performance improvements of the STGNNs rely heavily on complex model architectures (shao2022decoupled; lan2022dstagnn) to capture overall trends, yet low-predictability points in both temporal and spatial dimensions remain difficult to capture. Instead of further increasing model complexity by adding more weighted parameters, we propose to capture complex spatio-temporal dependencies through explicit memory storage and retrieval mechanisms, as demonstrated in Figure 2.
@@
To bridge the gap in applying RAG to STF and further address the aforementioned challenges, we propose RAST (Retrieval-Augmented Spatio-Temporal forecasting), a universal framework that integrates retrieval-augmented mechanisms with spatio-temporal modeling for traffic prediction and pre-trained model enhancements. Specifically, ❶ our approach maintains a vector-based dual-dimension spatio-temporal retrieval store. During training, our method decouples input data into spatial and temporal encodings used to update the store while constructing the context-aware queries through residual fusion. ❷ Specific retrievers then search and project dual-dimension retrieval embeddings under queries. ❸ Finally, we fuse decoupled retrieval embeddings with current queries through a cross-attention module and obtain final predictions via the universal backbone predictor. Our key contributions are summarized as follows:
@@
  • Comprehensive Empirical Validation: Extensive experiments on six real-world datasets demonstrate that our proposed method effectively captures complex spatio-temporal patterns while maintaining high computational efficiency, achieving up to 24.75% improvement in average MAE compared to RPMixer on the SD dataset.
@@ 2 Related Work
is a fundamental task in numerous application domains such as traffic management, urban planning, and environmental monitoring (lv2014traffic; jin2023spatio; bi2023accurate; jiang2021dl; wavets). Early deep learning methods combined CNNs and RNNs (shi2015convolutional; hochreiter1997long) but struggled with non-Euclidean traffic networks. The emergence of Spatio-temporal Graph Neural Networks (STGNNs) addressed this limitation by integrating Graph Neural Networks (GNNs) (kipf2016semi) with temporal models (li2017diffusion; zhang2017deep; yu2017spatio; bai2020adaptive; guo2019attention; shao2022decoupled; liang2018geoman; wu2020connecting; wu2019graph; zheng2020gman). Despite these advancements, the performance improvements of STGNNs have begun to plateau due to limited contextual capacity and reliance on increasingly complex architectures (wang2020traffic; jin2024survey). This stagnation has prompted research toward integrating pre-trained models and Large Language Models (LLMs) (zhou2024one; liu2023spatio; jin2024position; yuan2024unist; shao2022pre) to enhance predictive capabilities. However, these methods still struggle to capture spatio-temporal heterogeneity in large-scale scenarios, while universal solutions for contextual capacity limitations in spatio-temporal modeling remain largely unexplored.
@@
has emerged as a transformative paradigm for enhancing large language model performance, particularly in knowledge-intensive tasks (kandpal2023large; lewis2020retrieval; gao2023retrieval). Traditional LLMs struggle with tasks requiring vast amounts of factual knowledge or domain-specific expertise due to finite parametric memory and limited context (roberts2020much; petroni2019language). RAG addresses this limitation by integrating external retrieval mechanisms that enable models to dynamically access relevant information from large knowledge bases during inference (zhu2024rageval; izacard2021leveraging; borgeaud2022improving), improving open-domain question answering, fact-checking, and few-shot learning tasks (ram2023context; shi2023replug). While the retrieval-augmented mechanism has been extensively studied in natural language processing, its application to vast spatio-scenarios presents significant opportunities, particularly given the similar challenges of constrained model capacity.
3 Methodology
In this section, we present RAST, a Retrieval Augmented Spatio-Temporal forecasting framework as illustrated in Figure 3. Our approach consists of five core components working synergistically: 1) Decoupled Encoder Layers that convert raw inputs into vectorized embedding representations, 2) Query Generator that constructs fusion queries via residual fusion, 3) Retrieval Store that maintains and fine-grained historical patterns 4) ST-Retriever that retrieves embeddings based on queries and further fuses them with cross-attention blocks, and 5) Backbone Predictor that accommodates diverse pre-trained STGNNs or simple predictors (e.g. MLP) for enhanced forecasting performance.
3.1 Problem Formulation
@@
3.2 Data Encoding and Query Construction
@@
Following recent advances in spatio-temporal modeling (shao2022spatial), we employ decoupled encoder layers to separately process temporal and spatial information. This encoding module initially captures basic characteristics (e.g., cyclicity for temporal, regionality for spatial) formulated as follows:
@@
$\mathbf{E}_{f}^{(l+1)}$ $=\text{LayerNorm}(\mathbf{E}_{f}^{(l)}+\text{FFN}(\mathbf{E}_{f}^{(l)})),$ (5)
@@
3.3 Spatio-temporal Retrieval Store
@@
$\mathcal{M}_{sp}^{(i)}$ $=\{\mathbf{v}^{(i)}_{sp},\mathbf{m}_{sp}^{(i)}\}_{i=1}^{|\mathcal{M}_{sp}|}$ (7)
$\mathcal{M}_{tp}^{(j)}$ $=\{\mathbf{v}_{tp}^{(j)},\mathbf{m}_{tp}^{(j)}\}_{j=1}^{|\mathcal{M}_{tp}|}$ (8)
@@
To enable fast similarity search, our retrieval store utilizes the Facebook AI Similarity Search (FAISS) library (douze2024faiss; johnson2019billion) for efficient similarity-based indexing. Given history queries $\mathcal{Q}_{st}$ and current state $e$, we maintain and compute indices $\mathcal{I}$ for temporal and spatial embeddings (more details in Appendix D):
@@
Information-Theoretic ST-Retriever.
@@
$\mathcal{E}_{s}$ $=\text{Retriever}(\mathbf{E}_{sp},\mathcal{I}_{s},k)=\{(\mathbf{v}^{(i)}_{s},\omega^{(i)}_{s})\}_{i=1}^{k}$ (13)
@@
Momentum-Based Memory Management.
@@ 3.4 Prediction and Optimization
STDN 17.77 28.63 21.37 20.86 32.63 15.26 20.08 33.73 9.29 19.19 28.53 14.99
RAST (Ours) 15.20 25.81 16.12 18.39 29.93 12.43 19.52 32.73 8.23 14.16 23.33 9.27
Table 1: Performance comparison assessed by averaging over all 12 prediction steps with baseline models on the PEMS03, 04, 07, 08 datasets. Bold: best; Underline: second best.
4 Experiments
@@ 4.1 Experimental Settings
We conducted comprehensive experiments across six diverse traffic networks: 1) PEMS03, PEMS04, PEMS07, PEMS08 (song2020spatial) datasets, and 2) the large-scale dataset San Diego (SD), Greater Bay Area (GBA) as introduced in LargeST (liu2023largest). Detailed statistics of these datasets are given in Appendix B. We compare our RAST with 21 classic or advanced baselines, which are categorized into 3 groups: (i) Non-spatial methods: ARIMA (box2015time)VAR (lutkepohl2005new), SVR (awad2015support), LSTM (hochreiter1997long), TCN (lea2017temporal), Transformer (vaswani2017attention), NHiTS (challu2022nhitsneuralhierarchicalinterpolation), iTransformer (liu2023itransformer), TimeMixer (wang2024timemixer); (ii) GNN-based Spatial-temporal Models DCRNN (li2017diffusion), STGCN (yu2017spatio), ASTGCN (guo2019attention), GWNet (wu2019graph), LSGCN (huang2020lsgcn), STSGCN (song2020spatial), STFGNN (li2021spatial), STGODE (fang2021spatial), DSTAGNN (lan2022dstagnn), AGCRN (bai2020adaptive), D2STGNN (shao2022decoupled), and (iii) Other Enhanced Approaches: Z-GCNETs (chen2021multimodal), TAMP (chen2022tamp), STKD (wang2024spatial), RPMixer (yeh2024rpmixer), STDN (cao2025spatiotemporal). More details of the baselines are given in Appendix A.
@@
Performance is evaluated using standard metrics, including MAE, RMSE, and MAPE. We use 12 historical time steps to forecast the next 12 steps and calculate the average across horizons 3, 6, and 12. For the fairness of the experiment, we only utilized the simple MLP trained from scratch as the backbone predictor in the following experiments. All models are implemented in PyTorch and trained on Nvidia RTX A6000 GPUs with consistent hyperparameter tuning protocols to ensure fair comparison. The models are trained using the Adam optimizer with a learning rate of 0.002, a batch size of 32, and a maximum of 300 epochs, applying an early stopping strategy. More detailed experimental settings are provided in Appendix C.
Dataset Methods Horizon 3 Horizon 6 Horizon 12 Average
@@
RAST (Ours) 15.84 26.41 10.15 18.55 31.56 12.13 22.18 39.43 15.38 18.39 31.96 12.19
GBA LSTM 20.38 33.34 15.47 27.56 43.57 23.52 39.03 60.59 37.48 27.96 44.21 24.48
@@
RAST (Ours) 17.71 29.29 13.72 20.86 34.26 16.63 24.97 41.33 21.25 20.64 34.47 16.63
Table 2: Large-scale traffic forecasting performance comparison of our RAST and baselines on the SD and GBA datasets. Our RAST achieves the best performance across all prediction horizons and metrics. Bold: best; Underline: second best.
4.2 Performance Evaluation (RQ1)
Table 1 presents a comprehensive comparison of various approaches for traffic forecasting tasks across PEMS datasets. While the results on PEMS03 show room for improvement, RAST consistently outperforms state-of-the-art baseline methods on the remaining datasets. On the PEMS07 dataset, RAST achieves a MAE of 19.52, surpassing DSTAGNN by 8.87%. On the PEMS08 dataset, RAST reduces RMSE by 1.65 compared to competitive STKD, while on the PEMS04, it outperforms all baselines with an MAE of 18.39 and RMSE of 29.93.
Table 2 extends our evaluation to larger datasets, while the performance on the horizon 3, horizon 6, horizon 12, and the average of the whole 12 horizons are reported. Our method maintains its advantage on larger traffic networks, achieving the best results at each horizon. On the SD dataset, RAST achieves an average MAE of 18.39, RMSE of 31.96, and MAPE of 12.19%, representing improvements compared to the second-best method STGODE. For the long-term prediction (Horizon 12), our method achieves the best MAE of 22.18 and RMSE of 39.43, highlighting its robustness in modeling complex long-range dependencies. On even larger traffic datasets GBA, RAST still surpasses baselines across all metrics and horizons, outperforming strong baselines like DSTAGNN and RPMixer. These results demonstrate the scalability of RAST to model complex spatio-temporal dependencies in large-scale traffic networks.